# Toward Human-AI Alignment in Large-Scale Multi-Player Games

### v1

@@ 6 Discussion
In this work, we proposed a Task-sets-based framework for (i) understanding human behavior in large-scale multiplayer games, and (ii) assessing the alignment of AI agents with humans. Specifically, we apply this framework to examine the alignment of a proof of concept GPT-based AI agent trained to play Bleeding Edge. We propose a simultaneous affordance-completion analysis of these task-sets as a measure of human-agent similarity. We use this analysis to examine interpretable behavioral axes, allowing for richer comparisons than what policy differences alone allow for.
Our analysis of the human data shows that the task-sets capture meaningful factors in variation in human behavior along three behavioral axes: fight-flight, explore-exploit, and solo-multi-agent tendencies. However, while humans show substantial differences along these dimensions, data from our AI agent does not mirror the variations observed in human behavior. We take this as evidence that these AI agents are not aligned with humans, which suggests that cooperation between humans and these AI agents might be challenging. In particular, these agents may exhibit behaviors not easily anticipated by human collaborators, potentially leading to a less engaging and immersive gameplay experience for human players.
Broader impact: Our work shows that the alignment of transformer-based models trained on next token prediction may not always be inherent, and it may require specialized training techniques such as supervised fine-tuning (Gunel et al., 2020; Wortsman et al., 2022; Lee et al., 2022; Kirichenko et al., 2022) and Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020; Glaese et al., 2022; Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2023; Azar et al., 2023) to achieve alignment. This underscores the importance of our framework for measuring alignment in AI agents.
@@ A.1 Acknowledgements
We will add these after the review process to preserve anonymity.
A.2 Software and data
The source code for the model presented in this paper, a data sample, and the analysis code can be made available upon acceptance (for the camera-ready version).
A.3 Why Bleeding Edge
@@
A.4 AI Agent details
A.4.1 Training
@@
A.4.2 Observation Encoder
The model is trained on sequences of $T=128$ images where each image is reshaped to $128\times 128\times 3$ and then divided by 255 to ensure its value lies in range [0, 1]. A custom ResNet (He et al., 2016) with 18.6M parameters is used to embed each image observation ($O\in R^{3\times 128\times 128}$) independently into a vector. The first layer is a 2D convolutional network with kernels of shape $5\times 5$, a stride of 3, and a padding of 1 and maps to 64 channel dimension. We then apply GELU (Hendrycks & Gimpel, 2016) activation. This is followed by 4 ConvNext (Liu et al., 2022) and downsampling blocks. Each downsampling block applies group normalization and a convolution layer with kernel of shape $3\times 3$, stride of 2, and padding of 1, doubling the number of channels. We again apply GELU activation followed by another 2D convolutional network with a kernel of shape $3\times 3$, stride of $1\times 3$ and padding of 0. For each input image, the output of the encoder is a 1024D embedding.
A.5 Task-set definitions
A.5.1 Fight-Flight

### v2

@@ 6 Discussion
We propose a Task-sets-based framework for (i) understanding human behavior in large-scale multiplayer games, and (ii) assessing the alignment of AI agents with humans. We apply this framework to examine the alignment of a proof of concept GPT-based AI agent trained to play Bleeding Edge. Through simultaneous affordance-completion analysis of task-sets, we examine interpretable behavioral axes, allowing for richer comparisons than what policy differences alone allow for.
Our analysis of the human data shows that the task-sets capture meaningful factors of variation in human behavior along the three behavioral axes. While humans show substantial differences along these dimensions, data from our AI agent does not mirror the variations observed in human behavior. We take this as evidence that these AI agents are not aligned with humans. See A.7 for future work, broader impact and societal implications.
@@ A.1 Acknowledgements
We would like to acknowledge Tabish Rashid at Microsoft Research for support and guidance on the AI agent. We would also like to acknowledge the Ninja Theory team (that developed Bleeding Edge) for their feedback during the course of this project.
A.2 Why Bleeding Edge
@@
A.3 AI Agent details
A.3.1 Training
@@
A.3.2 Observation Encoder
The model is trained on sequences of $T=128$ images where each image is reshaped to $128\times 128\times 3$ and then divided by 255 to ensure its value lies in range [0, 1]. A custom ResNet He et al. (2016) with 18.6M parameters is used to embed each image observation ($O\in R^{3\times 128\times 128}$) independently into a vector. The first layer is a 2D convolutional network with kernels of shape $5\times 5$, a stride of 3, and a padding of 1 and maps to 64 channel dimension. We then apply GELU Hendrycks and Gimpel (2016) activation. This is followed by 4 ConvNext Liu et al. (2022) and downsampling blocks. Each downsampling block applies group normalization and a convolution layer with kernel of shape $3\times 3$, stride of 2, and padding of 1, doubling the number of channels. We again apply GELU activation followed by another 2D convolutional network with a kernel of shape $3\times 3$, stride of $1\times 3$ and padding of 0. For each input image, the output of the encoder is a 1024D embedding.
A.4 Task-set definitions
A.4.1 Fight-Flight
@@ A.5 Character types
A.6 AI Rollouts
For each game, we initialize the game at that state and let the AI agent play from that state onwards for a duration of $1$ minute. Except for the AI agent, all other players in the game are non-players who aren’t controlled and simply continue to repeat the latest action that the corresponding human player acted during the real human gameplay (e.g., if the player was moving forward when the game was originally recorded at that state, they will continue moving forward). They do not respond to the AI agent’s actions, however, their features get affected based on AI agent’s actions. For instance, their health decreases when the AI agent attacks them. We assume that each rollout is a different AI player given the stochasticity e.g., due to sampling the action from the output probability distribution.
Figure A.15: AI agent action accuracy curves over training. The Xbox controller action space consists of 12 discrete buttons and two joysticks. Top: Button accuracy curves for all the buttons (percentage of times the predicted value of all buttons matches the expected value) on the left; and for individual buttons (mean percentage of times the predicted value of each individual button matches its expected value). Middle: Left joystick accuracy curves for $x$ and $y$ components of the joystick. This joystick controls character movement. Bottom: Right joystick accuracy curves for $x$ and $y$ components of the joystick. This joystick controls the camera.
Figure A.16: AI agent decomposed loss curves over training. The Xbox controller action space consists of 12 discrete buttons and two joysticks with $x$ and $y$ components each of which is discretized into 11 bins. Top: Button loss curves for all 12 buttons. Middle: Left joystick loss curves for X and Y components of the joystick. This joystick controls character movement. Bottom: Right joystick loss curves for X and Y components of the joystick. This joystick controls the camera.
A.7 Future work, impact and implications
Broader impact: Our work shows that the alignment of transformer-based models trained on next token prediction may not always be inherent, and it may require specialized training techniques such as supervised fine-tuning Gunel et al. (2020); Wortsman et al. (2022); Lee et al. (2022); Kirichenko et al. (2022) and Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2017); Stiennon et al. (2020); Glaese et al. (2022); Ouyang et al. (2022); Bai et al. (2022); Rafailov et al. (2023); Azar et al. (2023) to achieve alignment. This underscores the importance of our framework for measuring alignment in AI agents.
Broader societal implications: Aligning AI with humans along fight-flight responses can help address ethical and moral questions about the use of AI in simulated (and potentially real) conflict situations, defense systems, and decision-making processes that involve risk, uncertainty, and potential harm to individuals and communities. AI alignment with human preferences for solo or multiplayer gameplay can influence social interactions, cultural norms, and community dynamics in gaming and entertainment, shaping how individuals and groups engage with AI-driven gaming experiences, collaborate with AI agents, and form online (and potentially offline) communities. Human-AI alignment along exploration or exploitation can foster innovation, creativity, and adaptive learning in various domains, including research, development, entrepreneurship, and education.
A.8 Ethics review program for human data collection
At Microsoft Research, we have an internal ethics review program that helped us work with human data in a way that ensured respect and protection of the rights of human participants contributing to our research.
Our institutional Review Board (IRB) approved all the data collection for the data used in this paper. For the player recordings, we received ethics approval (IRB 10601) from our organization’s Ethics Review Program. Our organization’s Institutional Review Board (IRB) has been officially registered with the U.S. Health and Human Services’ Office for Human Research Protections (OHRP) since 2017.
During data collection, human players playing Bleeding Edge are subjected to a pop up when they log in to the game for the first time to agree to the terms (end-user license agreement). The behavioral data collected does not contain any personally identifiable information.